Meeting ITk LPNHE

    • 10:30 10:50
      General news 20m



      • Module 084 bonded fully automatically. See report last Friday.
      • Module 016: problem for aligning due to glue… Also glue in the HV hole, difficult to wirebond but the team will try
      • Brigitte and Pascal to contact Julien to obtain parameters IREF TRIM Bit BondTable
      • Need to change the program due to different sensor thickness. Want to use module 7 to test. Brigitte will write an email to Francesco to know where module 7 is located
      • Reina will meet Brigitte, Pascal and Marc tomorrow to look at the database to obtain serial numbers


      • No major activities
      • Next step: no gluing until the dummy module is wire bonded. To be done when we are back at the end of August/beginning of September


      • Testing PC is not working… Activities stop until end of August. Members of the team on holidays
      • Luc placed the order for the plasma tube last week. Delay of 14 weeks, it will arrive ~October. Luc checking the setup with the Japanese team


      • Reina assembled (in the database) the 7 thin modules we received at LPNHE from Bonn
      • Mykola made progress with the scripts to include metrology information in the dB. Work will continue mid-August when he is back from holidays


      • Next meeting: August 3rd
      • For Yann, from Giovanni:  regarding the small aluminum slabs, we can work with something which is maximum 5% of large-z metals, not more. Thanks!